Operations Report – January 17th

Operations Report

Crew 219 Operations Report 17-JAN-2020

Sol: 12

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 13

From what time last night: 18:00.

This morning off at 07:00.

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 80%

9:00 = 98%; 10:00 = 99%; 11:00 = 100%; 13:00 = 95%; 15:00 = 87%; 17:00 = 80%;

Diesel Reading – 55%

Station Propane Reading- 30%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147842,6

Water (static tank) (gal) – 231

Static to loft Pump used –Yes

Water in Green Hab (gal): 142.85

Water in Science Dome (gal): 0

Toilet Tank Emptied –No

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- No

Starting Hours: 126.4

Beginning Charge: 100%

Ending Hours:

Currently Charging: Yes

Ending Charge:

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours: 132.4

Beginning Charge: 100%

Ending Hours:

Currently Charging: Yes

Ending Charge:

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used to retrieve Crew Car, on Campus
CrewCar used and why, where? Retrieved from Shop, on Campus

Journalist Report – January 17th

Crew 219 Journalist Report, 17 JAN 2020

Author: Nathan Hadland, Crew 219 Executive Officer

Sol 12

And That’s All Folks

We ended sim today at 1700 hours. We emerged from the Hab and collectively breathed in the fresh air and watched the sunset. The crew took in the beauty of the Utah desert without spacesuits for the first time in two weeks. I looked around at the crew and am content with the relationships I have strengthened and the work we have accomplished during our mission. Some of our research will be incorporated into senior theses, PhD dissertations, conference abstracts, and white papers. I believe that the impact of our work will extend far beyond this mission and will assist future MDRS crews as well as human space exploration as a whole. As the president of the Astrobiological Research and Education Society (ARES) at the Florida
Institute of Technology, my job is to ensure that the MDRS crews we organize reflect the overall values and vision of our organization. MDRS Crew 219 has accomplished that goal.

We woke up this morning to a hail storm and intense wind. For breakfast, Dave broke out his last can of bacon and prepared us a delicious breakfast to kick off the day of furious scrubbing and typing. Today was largely spent thoroughly cleaning the Hab and finishing up our mission summary. Dave (CO) and I started with a document of over 6,000 words and truncated it down to below the 2,000-word limit. That was an intense process! For lunch, Hannah made us cornbread and a compilation of the foods we had left. For dessert, we feasted on the leftover chocolate chip oatmeal cookies.

After ending sim, we went outside and began the tedious and brutal process of cleaning the mud-covered rovers in the cold using sponges and scrapers. After comms window tonight, we will enjoy the world’s best chai tea prepared by our geologist, Abdul as well as apple spice cake prepared by Hannah. We are also going to watch the MDRS essential film: The Martian. Overall, I had a phenomenal time at the Mars Desert Research Station, and I am excited for future ARES crews to return.

Greenhab Report – January 17th

Crew 219 GreenHab Report 17-JAN-20

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 24.3 °C

10:18 A.M.

Floor Unit: 21 °C

2:32 P.M.

Floor Unit: 20 °C

5:15 P.M.

Floor Unit: 20°C

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 9.18 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – 152.03 Gal.

Time(s) of watering for crops:

10:18 A.M.

SEASONAL PLANTING HISTORY:

Change to crops: Almost all of the crops look exceptionally bigger, brighter, and healthier

OTHER NOTES:

1st sprouts: The white onions that were planted a couple of days ago started to sprout, as is the green apple plant.

Harvest: At 2:32 P.M., 100 grams of swiss chard and 2 grams of chives were harvested.

Other: Since there is now only one thermometer that I am able to use within the GreenHab, I am only able to take floor temperature readings. I needed to constantly check on the temperature within the GreenHab to make sure that it was not too hot for the plants.

Sol Summary – January 17th

Crew 219 Sol Summary Report 17-01-2020

Sol: 12

Summary Title: Mission Complete.

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 12 started with gourmet coffee and a hearty breakfast of bacon and eggs. After breakfast D. Masaitis performed a walk-around inspection of the campus with Handbook in hand, to annotate any deficiencies that needed to be fixed before check-out, and passed the list to the crew. As the crew began touching up the modules and Hab, D. Masaitis sat down to truncate the rough draft of the crew’s Mission Summary from almost 7,000 words of eloquent academic prose. The crew had completed their cleaning tasks by lunchtime, and spent much of the afternoon working on the 2000 piece Mars puzzle. N. Hadland and H. Blackburn joined D. Masaitis in The Great Revision, and the Mission Summary was complete by 1630 hours.

Crew 219 ended Simulation at 1700 hours, and stepped out into the waning daylight to clean exterior areas, completing the work by sundown. We are proud of the work that we have accomplished here, and look forward to meeting Crew 220 tomorrow!

Look Ahead Plan: Meeting and Training Crew 220!

Anomalies in work: None

Weather: Chilly and sunny all day

Crew Physical Status: OUTSTANDING

EVA: NONE

Reports to be filed:

Sol Summary

Operations Report

GreenHab Report

Mission Summary (In lieu of Research Report)

Food Inventory

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested:

None